# Center for Advanced Intelligence Project, RIKEN 103-0027, Tokyo, Japan

### v1

@@ Abstract
By using the squared slack variables technique, we show that a general polynomial complementarity problem can be formulated as a system of polynomial equations. Thus, the solution set of such a problem is the image of a real algebraic set under a certain projection. This paper points out that, generically, this polynomial system has finitely many complex zeros. In such a case, we use techniques from symbolic computation to compute a univariate representation of the solution set. Consequently, univariate representations of special solutions, such as least-norm and sparse solutions, are obtained. After that, enumerating solutions boils down to solving problems governed by univariate polynomials. We also provide some experiments on small-scale problems with worst-case scenarios. At the end of the paper, we propose a method for computing approximate solutions to copositive polynomial complementarity problems that may have infinitely many solutions.
@@ 1 Introduction
where $\langle\alpha\mid\beta\rangle$ is the usual scalar product of two vectors $\alpha,\beta$ in the Euclidean space $\operatorname{\mathbb{R}}^{n}$. If $f$ is an affine mapping, the problem becomes a linear complementarity problem, first introduced by Cottle and Dantzig cottle1968complementary . If $f$ is a real polynomial mapping, i.e. all the components of $f$ are polynomial, then the problem is called a polynomial complementarity problem (PCP, for short) gowda2016polynomial . Moreover, PCP is a special case of a polynomial variational inequality hieu2020solution .
@@
We introduce a class of PCPs given by polynomial mappings $f$ such that the polynomial systems $\{x_{1}f_{1},\dots,x_{n}f_{n}\}$ have finitely many complex zeros. A PCP belonging to this class is called $\operatorname{\mathcal{D}}_{0}$, and it should be highlighted that this class is almost everywhere, see Proposition 2. Using Gröbner bases, the shape lemma, and results from computational commutative algebra, we prove in Theorem 4.1 the existence of a univariate representation for the solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. Based on the proof, we design an algorithm to compute such a representation, see Algorithm 1. Additionally, algorithms are proposed to compute univariate representations of least-norm or sparse solutions. Once a univariate representation is obtained, we can enumerate the solutions using numerical computation. Furthermore, we discuss the sharp upper bound on the number of solutions of $\operatorname{\mathcal{D}}_{0}$ problems based on the well-known Bézout bounds.
@@ 2.1 Polynomial complementarity problems
Let $f=(f_{1},\dots,f_{n}):\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ be a real polynomial mapping. The polynomial complementarity problem, introduced in gowda2016polynomial , for the function $f$ is the problem of finding a solution $\alpha\in\operatorname{\mathbb{R}}^{n}$ of the following system:
@@
$\min\{\|\alpha\|_{0}:\alpha\in\operatorname{Sol}(f)\},$ (2.5)
@@ 2.2 Zero-dimensional ideals and the Shape Lemma
Let $<$ be a monomial ordering on $\operatorname{\mathbb{R}}[x]$ and $\operatorname{\mathcal{I}}\neq\{0\}$ be an ideal of $\operatorname{\mathbb{R}}[x]$. We denote by $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$ the set of all leading terms $\operatorname{LT}_{<}(p)$ for $p\in\operatorname{\mathcal{I}}$, and by $\left\langle\operatorname{LT}_{<}(\operatorname{\mathcal{I}})\right\rangle$ the ideal generated by the elements of $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$. A subset $G=\{p_{1},\dots,p_{s}\}$ of $\operatorname{\mathcal{I}}$ is said to be a Gröbner basis of $\operatorname{\mathcal{I}}$ with respect to the monomial order $<$ if
@@
Note that a Gröbner basis of $\operatorname{\mathcal{I}}$ really is a generating set of the ideal $\operatorname{\mathcal{I}}$, i.e. $\left\langle p_{1},\dots,p_{s}\right\rangle=\operatorname{\mathcal{I}}$ and every ideal in $\operatorname{\mathbb{R}}[x]$ has a Gröbner basis (see (cox2013, , Cor. 6)). A Gröbner basis $G$ is reduced if the two following conditions hold: i) the leading coefficient of $p$ is $1$, for all $p\in G$; ii) for all $p\in G$, there are no monomials of $p$ lying in $\left\langle\operatorname{LT}_{<}(G\setminus\{p\})\right\rangle$. Every ideal $\operatorname{\mathcal{I}}$ has a unique reduced Gröbner basis. We refer the reader to cox2013 for more details.
Let $\operatorname{\mathcal{I}}$ be a zero-dimensional and radical ideal in $\operatorname{\mathbb{R}}[x]$, and $G$ be its reduced Gröbner basis with respect to the lexicographic ordering with $x_{1}<_{\operatorname{lex}}\cdots<_{\operatorname{lex}}x_{n}$. One says that $\operatorname{\mathcal{I}}$ is in shape position if $G$ has the following shape:
@@ 3 The squared slack variables technique for general PCPs
Furthermore, from the fact that #$\#\sqrt{c}$ is finite, for $c\in\operatorname{\mathbb{C}}$, we conclude that the set $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite if any only if $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$ so is.
@@
Consider a linear complementarity problem $\operatorname{CP}(f)$ in two variables, where $f(x_{1},x_{2})=(x_{2}-1,x_{1}-1)$. We use four slack variables $(z_{1},\dots,z_{4})$ and consider the following system of polynomials in $\operatorname{\mathbb{R}}[x_{1},x_{2},z_{1},\dots,z_{4}]$:
@@ 4 Existence of univariate representations for $\operatorname{\mathcal{D}}_{0}$ problems
Let $\operatorname{\mathbb{R}}_{d}[x]$ be the set of all polynomials with real coefficients of degree at most $d$. Recall that the dimension of $\operatorname{\mathbb{R}}_{d}[x]$, denoted by $\rho$, is finite and the linear space $\operatorname{\mathbb{R}}_{d}[x]^{n}$ is isomorphic to $\operatorname{\mathbb{R}}^{n\times\rho}$. For every $f=(f_{1},\dots,f_{n})$ in $\operatorname{\mathbb{R}}_{d}[x]^{n}$, $f_{i}$ can be written (uniquely) with coefficients $a_{i}\in\operatorname{\mathbb{R}},Q_{i}\in\operatorname{\mathbb{R}}^{\rho-1}$ as follows: $f_{i}(x)=a_{i}+Q_{i}X^{T}$, where
$X=(x_{1},\dots,x_{n},x_{1}^{2},x_{1}x_{2},\dots,x_{n}^{2},\dots,x_{n}^{d})^{T}.$ (4.1)
Recall that a property is said to be generic on $\operatorname{\mathbb{R}}_{d}[x]^{n}$ if it holds on a subset $\mathcal{O}\subset\operatorname{\mathbb{R}}_{d}[x]^{n}$ such that its complement $\operatorname{\mathbb{R}}_{d}[x]^{n}\setminus\mathcal{O}$ has Lebesgue measure zero.
The following proposition relates to the genericity of $\operatorname{\mathcal{D}}_{0}$ whose proof is postponed to the Appendix.
@@
The property of being $\operatorname{\mathcal{D}}_{0}$ is generic in $\operatorname{\mathbb{R}}_{d}[x]^{n}$.
@@ 4.1 A sharp bound on the cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$
It should be highlighted that if we use $m=2n$ squared slack variables as in Remark 3, the cardinality of the complex algebraic set can increase $2^{2n}$ times. However, in the setting of the complementarity problem, the multiplier is reduced from $2^{2n}$ to $2^{n}$, as in the following proposition. Recall that the degree of the polynomial mapping $f$ is the following number:
@@
Applying Bézout’s theorem (cox2013, , Chapter 8, §7) to the system defining $\operatorname{\mathcal{I}}[f]$, one has #$\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$. Therefore, the second inequality in (4.2) is obtained. ∎
@@
Let $d$ be fixed. Then the maximum number of solutions of $\operatorname{CP}(f)$, where $f$ is of degree $d$ and is $\operatorname{\mathcal{D}}_{0}$, does not exceed $(d+1)^{n}$. Indeed, it is clear that #$\#\operatorname{Sol}(f)\leq\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$, where the last number is the Bézout bound of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$.
Moreover, if $d\in\operatorname{\mathbb{N}}$ is even, then the maximum is reached, i.e., there exists $p\in\operatorname{\mathbb{R}}[x]^{n}$ of degree $d$ and being $\operatorname{\mathcal{D}}_{0}$ such that #$\#\operatorname{Sol}(p)=(d+1)^{n}$. We consider the polynomial map $p$ given by $p_{i}(x)=\prod_{j=1}^{d}(x_{i}-j),\ i\in[n]$. Clearly, $x_{i}p_{i}$ has $d+1$ non-negative (real) roots, listed by $\{0,1,\dots,d\}$. It is not difficult to check that

### v3

@@ Abstract
By using the squared slack variables technique, we demonstrate that the solution set of a general polynomial complementarity problem is the image, under a specific projection, of the set of real zeroes of a system of polynomials. This paper points out that, generically, this polynomial system has finitely many complex zeroes. In such a case, we use symbolic computation techniques to compute a univariate representation of the solution set. Consequently, univariate representations of special solutions, such as least-norm and sparse solutions, are obtained. After that, enumerating solutions boils down to solving problems governed by univariate polynomials. We also provide some experiments on small-scale problems with worst-case scenarios. At the end of the paper, we propose a method for computing approximate solutions to copositive polynomial complementarity problems that may have infinitely many solutions.
@@ 1 Introduction
where $\langle\alpha\mid\beta\rangle$ is the usual scalar product of two vectors $\alpha,\beta$ in the Euclidean space $\operatorname{\mathbb{R}}^{n}$. If $f$ is an affine mapping, the problem becomes a linear complementarity problem, first introduced by Cottle and Dantzig cottle1968complementary . If $f$ is a real polynomial mapping, i.e. all the components of $f$ are polynomial, then the problem is called a polynomial complementarity problem (PCP, for short) gowda2016polynomial . Moreover, a PCP is a special case of the polynomial variational inequality hieu2020solution .
@@
We introduce a class of PCPs given by polynomial mappings $f$ such that the polynomial systems $\{x_{1}f_{1},\dots,x_{n}f_{n}\}$ have finitely many complex zeros. A PCP belonging to this class is called $\operatorname{\mathcal{D}}_{0}$, and it should be highlighted that, as shown in Proposition 2, the property of being $\operatorname{\mathcal{D}}_{0}$ is generic. Using Gröbner bases, the Shape Lemma, and results from computational commutative algebra, we prove in Theorem 4.2 the existence of a univariate representation for the solution set of a $\operatorname{\mathcal{D}}_{0}$ problem. Based on the proof, we design an algorithm to compute such a representation, see Algorithm 1. Additionally, algorithms are proposed to compute univariate representations of least-norm or sparse solutions. Once a univariate representation is obtained, we can enumerate the solutions by using numerical computation. Furthermore, we discuss the sharp upper bound on the number of solutions of a $\operatorname{\mathcal{D}}_{0}$ problem based on the well-known Bézout bound.
@@ 2.1 Polynomial Complementarity Problems
Let $f=(f_{1},\dots,f_{n}):\operatorname{\mathbb{R}}^{n}\to\operatorname{\mathbb{R}}^{n}$ be a real polynomial mapping. The polynomial complementarity problem, introduced in gowda2016polynomial , for $f$ is the problem of finding a solution $\alpha\in\operatorname{\mathbb{R}}^{n}$ of the following system:
@@
$\min\{\|\alpha\|_{0}:\alpha\in\operatorname{Sol}(f)\},$
@@ 2.2 Zero-dimensional Ideals and the Shape Lemma
Let $<$ be a monomial ordering on $\operatorname{\mathbb{R}}[x]$ and $\operatorname{\mathcal{I}}\neq\{0\}$ be an ideal of $\operatorname{\mathbb{R}}[x]$. We denote by $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$ the set of all leading terms $\operatorname{LT}_{<}(p)$ for $p\in\operatorname{\mathcal{I}}$, and by $\left\langle\operatorname{LT}_{<}(\operatorname{\mathcal{I}})\right\rangle$ the ideal generated by the elements of $\operatorname{LT}_{<}(\operatorname{\mathcal{I}})$. A subset $G=\{p_{1},\dots,p_{s}\}$ of $\operatorname{\mathcal{I}}$ is said to be a Gröbner basis of $\operatorname{\mathcal{I}}$ with respect to the monomial ordering $<$ if
@@
Note that a Gröbner basis of $\operatorname{\mathcal{I}}$ really is a generating set of the ideal $\operatorname{\mathcal{I}}$, i.e. $\left\langle p_{1},\dots,p_{s}\right\rangle=\operatorname{\mathcal{I}}$ and every ideal in $\operatorname{\mathbb{R}}[x]$ has a Gröbner basis. A Gröbner basis $G$ is reduced if the two following conditions hold: i) the leading coefficient of $p$ is $1$, for all $p\in G$; ii) for all $p\in G$, there are no monomials of $p$ lying in $\left\langle\operatorname{LT}_{<}(G\setminus\{p\})\right\rangle$. Every ideal $\operatorname{\mathcal{I}}$ has a unique reduced Gröbner basis. We refer the reader to cox2013 for more details.
Let $\operatorname{\mathcal{I}}$ be a zero-dimensional and radical ideal in $\operatorname{\mathbb{R}}[x]$, and $G$ be its reduced Gröbner basis with respect to a lexicographic ordering with $x_{1}<_{\operatorname{lex}}\cdots<_{\operatorname{lex}}x_{n}$. One says that $\operatorname{\mathcal{I}}$ is in shape position if $G$ has the following shape:
@@ 3 The Squared Slack Variables Technique for General PCPs
Furthermore, from the fact that #$\#\sqrt{c}$ is finite, for $c\in\operatorname{\mathbb{C}}$, we conclude that the set $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$ is finite if any only if so is $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$.
@@
Consider a linear complementarity problem $\operatorname{CP}(f)$ in two variables, where $f(x_{1},x_{2})=(x_{2}-1,x_{1}-1)$. We use four slack variables $(z_{1},\dots,z_{4})$ and consider the following system of polynomials in $\operatorname{\mathbb{R}}[x_{1},x_{2},z_{1},z_{2},z_{3},z_{4}]$:
@@ 4 Existence of Univariate Representations for $\operatorname{\mathcal{D}}_{0}$ Problems
Let $\operatorname{\mathbb{R}}_{d}[x]$ be the set of all polynomials with real coefficients of degree at most $d$. Recall that the dimension of $\operatorname{\mathbb{R}}_{d}[x]$, denoted by $\rho$, is finite and the linear space $\operatorname{\mathbb{R}}_{d}[x]^{n}$ is isomorphic to $\operatorname{\mathbb{R}}^{n\times\rho}$. For every $f=(f_{1},\dots,f_{n})$ in $\operatorname{\mathbb{R}}_{d}[x]^{n}$, $f_{i}$ can be written (uniquely) with coefficients $a_{i}\in\operatorname{\mathbb{R}},\ Q_{i}\in\operatorname{\mathbb{R}}^{\rho-1}$ as follows: $f_{i}(x)=a_{i}+Q_{i}^{T}\operatorname{vec}(x)$, where
$\operatorname{vec}(x)=(x_{1},\dots,x_{n},x_{1}^{2},x_{1}x_{2},\dots,x_{n}^{2},\dots,x_{n}^{d})^{T}.$ (4.1)
Then, $f$ is written concisely $f(x)=a+Q^{T}\operatorname{vec}(x)$, where $a=(a_{1},\dots,a_{n})$ and $Q=(Q_{1},\dots,Q_{n})$.
4.1 Genericity of $\operatorname{\mathcal{D}}_{0}$ Problems
A subset $\mathcal{O}\subset\operatorname{\mathbb{R}}^{m}$ is said to be generic on $\operatorname{\mathbb{R}}^{m}$ if its complement $\operatorname{\mathbb{R}}^{m}\setminus\mathcal{O}$ has Lebesgue measure zero. Recall from (pham2016genericity, , Proposition 1.5) that if $\mathcal{O}\subset\operatorname{\mathbb{R}}^{m}$ is an open and dense semi-algebraic subset, then $\mathcal{O}$ is generic in $\operatorname{\mathbb{R}}^{m}$.
This subsection aims at proving that the property of being $\operatorname{\mathcal{D}}_{0}$ is generic, i.e. there exists a generic subset $\mathcal{O}$ in the parameter space $\operatorname{\mathbb{R}}^{n\times\rho}$ such that $a+Q^{T}\operatorname{vec}(x)$ is $\operatorname{\mathcal{D}}_{0}$ for every $(a,Q)\in\mathcal{O}$.
The following theorem (Theorem 4.1) is a version of Sard’s theorem with real parameters. Other versions can be found, for example, in (pham2016genericity, , Theorem 1.10) for the semi-algebraic setting (with real parameters), or in (safey2017, , Proposition B.3) for the algebraic setting (with complex parameters). The claim is considered folklore; however, a suitable reference for it has yet to be found. The proof is postponed to the appendix. The notions of real/complex manifolds and regular values can be found in monographs, for example, lee2003smooth ; lee2024complex .
Theorem 4.1
Let $V\subset\operatorname{\mathbb{C}}^{n}$ be a smooth manifold, and $\Phi:\operatorname{\mathbb{R}}^{m}\times V\to\operatorname{\mathbb{C}}^{n}$ be a polynomial mapping. Assume that $V$ is the difference of two algebraic sets and $0\in\operatorname{\mathbb{C}}^{n}$ is a regular value of $\Phi$. Then, there exists an open and dense semi-algebraic set $\mathcal{O}$ in $\operatorname{\mathbb{R}}^{m}$ such that 00 is a regular value of the mapping
$\Phi_{c}:V\to\operatorname{\mathbb{C}}^{n},\ \Phi_{c}(x)=\Phi(c,x),$
for any $c\in\mathcal{O}$.
@@ 4.1 Genericity of $\operatorname{\mathcal{D}}_{0}$ Problems
There exists an open and dense semi-algebraic subset $\mathcal{O}$ in $\operatorname{\mathbb{R}}^{n\times\rho}$ such that $f(x)=a+Q^{T}\operatorname{vec}(x)$ is $\operatorname{\mathcal{D}}_{0}$ for every $(a,Q)\in\mathcal{O}$.
Proof
Let $\ell$ be a subset of $[n]$ and $\mathcal{V}_{\ell}$ be the set of all $\alpha\in\operatorname{\mathbb{C}}^{n}$ such that $\alpha_{i}\neq 0$ if $i\in\ell$ and $\alpha_{i}=0$ if $i\in\bar{\ell}=[n]\setminus\ell$. Clearly, $\mathcal{V}_{\emptyset}=\{0\}$ and $\operatorname{\mathbb{C}}^{n}$ is the disjoint union of $\mathcal{V}_{\ell}$ for $\ell\subseteq[n]$.
Let $\ell\neq\emptyset$ be given. Without loss of generality, we assume that =$\ell=\{1,\dots,l\}$. For convenience, we write $x_{\ell}=(x_{1},\dots,x_{l})$ and 0=($0=(0_{\ell},0_{\bar{\ell}})$, where $0_{\bar{\ell}}$ is the zero $(n-l)$-vector. Then, the set $\mathcal{V}_{\ell}$ can be considered as the difference of two algebraic sets:
$\mathcal{V}_{\ell}=\operatorname{\mathbb{C}}^{l}\times\{0_{\bar{\ell}}\}\setminus\{(x_{\ell},0_{\bar{\ell}}):x_{1}\cdots x_{l}=0\}.$
We consider the polynomial mapping
$\Phi_{\ell}:\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n\times(\rho-1)}\times\operatorname{\mathbb{C}}^{l}\times\{0_{\bar{\ell}}\}\to\operatorname{\mathbb{C}}^{l}$
defined by
$\Phi_{\ell}(a,Q,x_{\ell},0_{\bar{\ell}})=\Big{(}a_{1}x_{1}+x_{1}Q_{1}^{T}\operatorname{vec}(x_{\ell},0_{\bar{\ell}}),\ \dots,\ a_{l}x_{l}+x_{l}Q_{l}^{T}\operatorname{vec}(x_{\ell},0_{\bar{\ell}})\Big{)},$
where $Q\in\operatorname{\mathbb{R}}^{(\rho-1)\times n}$ and $\operatorname{vec}(x)$ is given in (4.1) with $x=(x_{\ell},0_{\bar{\ell}})$. One sees that the restriction
$\Phi_{\ell}:\operatorname{\mathbb{R}}^{n}\times\operatorname{\mathbb{R}}^{n\times(\rho-1)}\times\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{l}$
is differentiable. Obviously, its Jacobian matrix with respect to $(a,Q,x_{\ell})$, denoted by $J_{\ell}$, has size $l\times(n\rho+l)$. Moreover, $J_{\ell}$ has a has the diagonal submatrix $\partial_{a}\Phi_{\ell}$, whose entries are the components of $x_{\ell}$. Since all coordinates of $x_{\ell}$ are non-zero, we conclude that $J_{\ell}$ is full rank. It follows that the origin $0_{\ell}$ in $\operatorname{\mathbb{C}}^{l}$ is a regular value of $\Phi_{\ell}$. By Theorem 4.1, there exists a generic set $\mathcal{O}_{\ell}$ in $\operatorname{\mathbb{R}}^{n\times\rho}$ such that for every $(a,Q)$ in $\mathcal{O}_{\ell}$, $0_{\ell}$ is a regular value of $\Phi_{\ell,a,Q}:\mathcal{V}_{\ell}\to\operatorname{\mathbb{C}}^{l}$ defined by $\Phi_{\ell,a,Q}(x_{\ell},0_{\bar{\ell}})=\Phi_{\ell}(a,Q,x_{\ell},0_{\bar{\ell}})$. It is clear that
$\Phi_{\ell,a,Q}^{-1}(0_{\ell})=\Big{\{}(x_{\ell},0_{\bar{\ell}})\in\mathcal{V}_{\ell}:a_{i}x_{i}+x_{i}Q_{i}^{T}\operatorname{vec}(x_{\ell},0_{\bar{\ell}})=0,i\in\ell\Big{\}},$
and according to the Regular Level Set Theorem (lee2024complex, , Chapter 2), this level set is a complex sub-manifold whose dimension is 00 or it is empty. Therefore, this set has finitely many point.
Finally, we observe that
$V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\subseteq\big{\{}0\big{\}}\cup\Big{(}\bigcup_{\ell\subseteq[n],\ell\neq\emptyset}\Phi_{\ell,a,Q}^{-1}(0_{\ell})\Big{)},$
@@ 4.2 A Sharp Bound on the Cardinality of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f,z])$
It should be highlighted that if we use $m=2n$ squared slack variables as in Remark 4, the cardinality of the complex algebraic set can increase up to $2^{2n}$ times the original. However, in the setting of the complementarity problem, the last number is reduced from $2^{2n}$ to $2^{n}$, as shown in the following proposition. Recall that the degree of the polynomial mapping $f$ is the following number:
@@
Applying Bézout’s theorem (cox2013, , Chapter 8, §7) to the system defining $\operatorname{\mathcal{I}}[f]$, one has #$\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$. Therefore, the second inequality in (4.2) holds. ∎
@@
Let $d$ be fixed. Then the maximum number of solutions of $\operatorname{CP}(f)$, where $f$ is of degree $d$ and is $\operatorname{\mathcal{D}}_{0}$, does not exceed $(d+1)^{n}$. Indeed, it is clear that #$\#\operatorname{Sol}(f)\leq\#V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])\leq(d+1)^{n}$, where the last number is the Bézout bound of $V_{\operatorname{\mathbb{C}}}(\operatorname{\mathcal{I}}[f])$. The maximum is reached, i.e., there exists $p\in\operatorname{\mathbb{R}}[x]^{n}$ of degree $d$ and being $\operatorname{\mathcal{D}}_{0}$ such that #$\#\operatorname{Sol}(p)=(d+1)^{n}$. We consider the polynomial map $p$ given by
${\color[rgb]{0,0,0}{p_{i}(x)=(-1)^{d}\prod_{j=1}^{d}(x_{i}-j),\ i\in[n]}}.$
Clearly, $x_{i}p_{i}$ has $d+1$ non-negative (real) roots, listed by $\{0,1,\dots,d\}$. It is not difficult to check that